OpenStudy (anonymous):

An airplane flew 3.5 hours with a 40 mph head wind. The return trip with a tail wind of the same speed took 2.5 hours. Find the speed of the plane in still air. 240 mph 220 mph 200 mph 180 mph

OpenStudy (anonymous):

let speed in still air be x mph .. speed with head wind = x-40 speed with tail wind = x+40 .. Distance = speed * time .. Headwind distance = 3.5(x-40) Tailwind distance = 2.5(x+40) .. 3.5(x-40) = 2.5(x+40) 3.5x-140=2.5x+100 x=140+100 x=240 mph speed of plane in still air ..